# DACAD: Domain Adaptation Contrastive Learning for Anomaly Detection in Multivariate Time Series

### v1

@@ Abstract
Time series anomaly detection (TAD) faces a significant challenge due to the scarcity of labelled data, which hinders the development of accurate detection models. Unsupervised domain adaptation (UDA) addresses this challenge by leveraging a labelled dataset from a related domain to detect anomalies in a target dataset. Existing domain adaptation techniques assume that the number of anomalous classes does not change between the source and target domains. In this paper, we propose a novel Domain Adaptation Contrastive learning for Anomaly Detection in multivariate time series (DACAD) model to address this issue by combining UDA and contrastive representation learning. DACAD’s approach includes an anomaly injection mechanism that introduces various types of synthetic anomalies, enhancing the model’s ability to generalise across unseen anomalous classes in different domains. This method significantly broadens the model’s adaptability and robustness. Additionally, we propose a supervised contrastive loss for the source domain and a self-supervised contrastive triplet loss for the target domain, improving comprehensive feature representation learning and extraction of domain-invariant features. Finally, an effective Centre-based Entropy Classifier (CEC) is proposed specifically for anomaly detection, facilitating accurate learning of normal boundaries in the source domain. Our extensive evaluation across multiple real-world datasets against leading models in time series anomaly detection and UDA underscores DACAD’s effectiveness. The results validate DACAD’s superiority in transferring knowledge across domains and its potential to mitigate the challenge of limited labelled data in time series anomaly detection.
1 Introduction
@@
Unsupervised Domain Adaptation (UDA) is a technique used to transfer knowledge from a labelled source domain to an unlabeled target domain, particularly useful when labelled data in the target domain are scarce or unavailable. Deep learning methods have become the predominant approach in UDA, offering advanced capabilities and significantly improved performance compared to traditional techniques. UDA’s approach is crucial in situations where deep models’ performance drops significantly due to the discrepancy between the data distributions in the source and target domains, a phenomenon known as domain shift Liu et al. (2022).
@@
Applications of UDA are diverse, ranging from image and video analysis to natural language processing and time series data analysis. In time series data analysis and anomaly detection, UDA can be particularly challenging, given that often i) the nature of time series data across different domains is complex and varied Wilson and Cook (2020); Liu et al. (2022), and ii) the number of anomalous classes changes between the source and target domains.
@@
For time series data specifically, UDA methods often employ neural architectures as feature extractors. These models are designed to handle domain adaptation, primarily target time series regression and classification problems Purushotham et al. (2016); Cai et al. (2021); Ozyurt et al. (2023); Wilson et al. (2020), focusing on aligning the major distributions of two domains. This approach may lead to negative transfer effects Zhang et al. (2022) on minority distributions, which is a critical concern in time series anomaly detection (TAD). The minority distributions, often representing rare events or anomalies, may be overshadowed or incorrectly adapted due to the model’s focus on aligning dominant data distributions, potentially leading to a higher rate of false negatives in anomaly detection. Since transferring knowledge of anomalies requires aligning minority distributions and the anomaly label spaces often have limited similarity across domains, existing methods face limitations in addressing anomaly detection in time series data. This highlights a significant gap in current UDA methodologies, pointing towards the need for novel approaches that can effectively address the unique requirements of anomaly detection in time series data.
@@
Furthermore, in the realm of time series anomaly detection, in the most recent model called ContextDA Lai et al. (2023), the discriminator aligns source/target domain windows without leveraging label information in the source domain, which makes the alignment ineffective. Our model leverages source labels and anomaly injection for better feature extraction, enhancing the alignment of normal samples. This is particularly vital due to the challenges in aligning anomalies across domains with different anomalous classes.
@@
In our study, we introduce Domain Adaptation Contrastive learning model for Anomaly Detection in time series (DACAD), a unique framework for UDA in multivariate time series anomaly detection, leveraging contrastive learning (CL). DACAD focuses on contextual representations. It forms positive pairs based on proximity and negative pairs using anomaly injection, subsequently learning their embeddings with both supervised CL in the source domain and self-supervised CL in the target domain. To sum up, we make the following contributions:
@@
  • We introduce DACAD, a pioneering CL framework for multivariate time series anomaly detection with UDA. It is a novel deep UDA model based on contrastive representation utilising a labelled source dataset and synthetic anomaly injection to overcome the lack of labelled data in TAD, distinguishing between normal and anomalous patterns effectively.
@@
  • Extending on the deep one-class classifier (DeepSVDD) Ruff et al. (2018), our proposed one-class classifier, the CEC, operates on the principle of spatial separation in the feature space by leveraging the existing label information in the source domain, aiming to bring “normal“ sample representations closer to the centre and distancing anomalous ones. This spatial adjustment is measured using a distance metric, establishing the basis for anomaly detection.
@@
  • Our comprehensive evaluation with real-world datasets highlights DACAD’s efficiency. In comparison with the recent TAD deep models and UDA models for time series classification and UDA for anomaly detection, DACAD demonstrates superior performance, emphasising the considerable importance of our study.
2 Related Works
@@
In this section, we provide a brief survey of the existing literature that intersects with our research. We concentrate on two areas: UDA for time series and deep learning for TAD.
@@
UDA for time series: In the realm of UDA for time series analysis, the combination of domain adaptation techniques with the unique properties of time series data presents both challenges and opportunities. Traditional approaches such as Maximum Mean Discrepancy (MMD) Long et al. (2018) and adversarial learning frameworks Tzeng et al. (2017) are geared towards minimizing domain discrepancies by developing domain-invariant features. These methods are vital in applications spanning medical, industrial, and speech data, with significant advancements in areas like sleep classification Zhao et al. (2021), arrhythmia classification Wang et al. (2021), and various forms of anomaly detection Lai et al. (2023), fault diagnosis Lu et al. (2021), and lifetime prediction Ragab et al. (2020).
@@
Time series specific UDA approaches like variational recurrent adversarial deep domain adaptation (VRADA) by Purushotham et al., pioneered UDA for multivariate time series, utilising adversarial learning with an LSTM networks Hochreiter and Schmidhuber (1997) and variational RNN Chung et al. (2015) feature extractor. Convolutional deep domain adaptation for time series (CoDATS) by Wilson et al., built on VRADA’s adversarial training but employed a convolutional neural network as the feature extractor. A metric-based method, time series sparse associative structure alignment (TS-SASA)  Cai et al. (2021) aligns intra and inter-variable attention mechanisms between domains using MMD. Adding to these, the CL for UDA of time series (CLUDA) Ozyurt et al. (2023) model offers a novel approach, enhancing domain adaptation capabilities in time series. All these methods share a common objective of aligning features across domains, each contributing unique strategies to the complex challenge of domain adaptation in time series classification data. However, they are ineffective when applied to TAD tasks. Additionally, ContextDA, introduced by Lai et al., is a TAD model that applies deep reinforcement learning to optimise domain adaptation, framing context sampling as a Markov decision process. However, it is ineffective when the anomaly classes change between source and target.
@@
Deep Learning for TAD: The field of Time Series Anomaly Detection (TAD) has advanced significantly, embracing a variety of methods ranging from basic statistical approaches to sophisticated deep learning techniques Schmidl et al. (2022); Audibert et al. (2022). Notably, deep learning has emerged as a promising approach due to its autonomous feature extraction capabilities Darban et al. (2022). TAD primarily focuses on unsupervised Yue et al. (2022); Hundman et al. (2018); Audibert et al. (2020); Xu et al. (2022) and semi-supervised methods Niu et al. (2020); Park et al. (2018) to tackle the challenge of limited labelled data availability. Unsupervised methods like OmniAnomaly Su et al. (2019), GDN Deng et al. (2021), and BeatGAN Zhou et al. (2019) are especially valuable in scenarios with sparse anomaly labels, whereas semi-supervised methods leverage available labels effectively.
@@
Advanced models such as LSTM-NDT Hundman et al. (2018) and THOC Shen et al. (2020) excel in minimizing forecasting errors and capturing temporal dependencies. However, models like USAD Audibert et al. (2020) face challenges with long time series data due to error accumulation in decoding. Additionally, unsupervised representation learning, exemplified by SPIRAL Lei et al. (2019) and TST Zerveas et al. (2021), shows impressive performance, albeit with scalability issues in long time series. Newer models like TNC Tonekaboni et al. (2021) and T-Loss Franceschi et al. (2019) aim to overcome these challenges using methods such as time-based negative sampling.
@@
Furthermore, contrastive representation learning, crucial in TAD for pattern recognition, groups similar samples together while distancing dissimilar ones. It has been effectively employed in TS2Vec Yue et al. (2022) for multi-level semantic representation and in DCdetector Yang et al. (2023), which uses a dual attention asymmetric design for permutation invariant representations.
3 DACAD
@@
Problem formulation: Given an unlabeled time series dataset $T$ (target), the problem is to detect anomalies in $T$ using a labelled time series dataset $S$ (source) from a related domain.
@@
In this section, we present DACAD, which ingeniously uses temporal correlations in time series data and adapts to differences between source and target domains. It starts with a labelled dataset in the source domain, featuring both normal and anomalous instances. Both real and synthetic (injected) anomalies aid in domain adaptation, ensuring training and improving generalisability on a wide range of anomaly classes.
@@
DACAD’s core is CL as inspired by a recent UDA for time series work Ozyurt et al. (2023), which strengthens its ability to handle changes between domains by improving the feature representation learning in source and target domains. As described in subsection 3.3, in the target domain, we use a self-supervised contrastive loss Schroff et al. (2015) by forming triplets to minimise the distance between similar samples and maximise the distance between different samples. Additionally, in the source domain, we leverage label information of both anomaly and normal classes and propose an effective supervised contrastive loss Khosla et al. (2020) named supervised mean-margin contrastive loss. A Temporal Convolutional Network (TCN) in DACAD captures temporal dependencies, generating domain-invariant features. A discriminator ensures these features are domain-agnostic, leading to consistent anomaly detection across domains. DACAD’s strength lies in its dual ability to adapt to new data distributions and to distinguish between normal and anomalous patterns effectively. In DACAD, time series data is split into overlapping windows of size $WS$ with a stride of 1, forming detailed source ($S$) and target ($T$) datasets. Windows in the source domain are classified based on anomaly presence: those with anomalies are marked as anomalous windows ($S_{\text{anom}}$), and others as normal ($S_{\text{norm}}$). Figure 1 shows the DACAD architecture, and Algorithm 1 details its steps. The following subsections explore DACAD’s components and their functionalities.
@@
5:Create $S_{\text{inj}}$, $T_{\text{inj}}$triangleright Anomaly Injection (3.1)
6:Form $S_{\text{triplets}}$ and $T_{\text{triplets}}$triangleright Pair Selection (3.2)
@@
8:    Compute $\phi^{R}(S)$,$\phi^{R}(T)$,$\phi^{R}(S_{\text{Triplets}})$,$\phi^{R}(T_{\text{Triplets}})$triangleright (3.3)
@@
11:    Compute $\mathcal{L}_{\text{Disc}}$ using Eq. (3), $\phi^{R}(S)$ and $\phi^{R}(T)$triangleright (3.4)
12:    Compute $\mathcal{L}_{\text{Cls}}$ using Eq. (4) and $\phi^{R}(S)$triangleright (3.5)
@@
3.1 Anomaly Injection
In the anomaly injection phase, we augment the original time series windows through a process of negative augmentation, thereby generating synthetic anomalous time series windows. This step is applied to all windows in the target domain ($T$) and all normal windows in the source domain ($S_{\text{norm}}$). We employ the anomaly injection method outlined in Darban et al. (2023), which encompasses five distinct types of anomalies: Global, Seasonal, Trend, Shapelet, and Contextual. This procedure results in the creation of two new sets of time series windows. The first set, $S_{\text{inj}}$, consists of anomaly-injected windows derived from the normal samples in the source domain. The second set, $T_{\text{inj}}$, comprises anomaly-injected windows originating from the target domain data.
3.2 Pair Selection
@@
3.3 Representation Layer ($\phi^{R}$)
In our model, we employ a TCN Lea et al. (2016) for the representation layer, which is adept at handling both multivariate time series windows. This choice is motivated by the TCN’s ability to capture temporal dependencies effectively, a critical aspect in time series analysis. The inputs to the TCN are the datasets and triplets from both domains, specifically $S$, $T$, $S_{\text{Triplets}}$, and $T_{\text{Triplets}}$. The outputs of the TCN, representing the transformed feature space, are

### v2

@@ Abstract
In time series anomaly detection (TSAD), the scarcity of labeled data poses a challenge to the development of accurate models. Unsupervised domain adaptation (UDA) offers a solution by leveraging labeled data from a related domain to detect anomalies in an unlabeled target domain. However, existing UDA methods assume consistent anomalous classes across domains. To address this limitation, we propose a novel Domain Adaptation Contrastive learning model for Anomaly Detection in multivariate time series (DACAD), combining UDA with contrastive learning. DACAD utilizes an anomaly injection mechanism that enhances generalization across unseen anomalous classes, improving adaptability and robustness. Additionally, our model employs supervised contrastive loss for the source domain and self-supervised contrastive triplet loss for the target domain, ensuring comprehensive feature representation learning and domain-invariant feature extraction. Finally, an effective Centre-based Entropy Classifier (CEC) accurately learns normal boundaries in the source domain. Extensive evaluations on multiple real-world datasets and a synthetic dataset highlight DACAD’s superior performance in transferring knowledge across domains and mitigating the challenge of limited labeled data in TSAD.
Index Terms:
Anomaly Detection, Time Series, Deep Learning, Contrastive Learning, Self-Supervised Learning, Domain Adaptation
I Introduction
@@
Unsupervised Domain Adaptation (UDA) is a technique used to transfer knowledge from a labelled source domain to an unlabeled target domain. Such an approach is particularly useful when labelled data in the target domain are scarce or unavailable [1]. Deep learning methods have become the predominant approach in UDA, offering advanced capabilities and significantly improved performance compared to traditional techniques. UDA is crucial in situations where the performance of deep models drops significantly due to the discrepancy between the data distributions in the source and target domains, a phenomenon known as domain shift [2].
@@
Applications of UDA are diverse, ranging from image and video analysis to natural language processing and time series data analysis. However, the time series domain differs greatly from the image and text data domains for which UDA methods are well developed [3, 4]. UDA can be particularly challenging for time series analysis, given that often i) the aspects of time series that may be relevant to different tasks can vary greatly, such as frequency, magnitude, rate of change in either of these, global shape or local shape [5, 2, 6], and ii) the number of anomalous classes changes between the source and target domains.
@@
For time series data specifically, UDA methods often employ neural architectures as feature extractors. These models are designed to handle domain adaptation, primarily target time series regression and classification problems [7, 8, 9, 10], focusing on aligning the major distributions of two domains. This approach may lead to negative transfer effects [11] on minority distributions, which is a critical concern in time series anomaly detection (TSAD). The minority distributions, often representing rare events or anomalies, may be overshadowed or incorrectly adapted due to the model’s focus on aligning dominant data distributions, potentially leading to a higher rate of false negatives in anomaly detection [12]. Since transferring knowledge of anomalies requires aligning minority distributions, anomaly label spaces often have limited similarity across domains, and existing methods face limitations in addressing anomaly detection in time series data. This highlights a significant gap in current UDA methodologies, pointing towards the need for novel approaches that can effectively address the unique requirements of anomaly detection in time series data.
@@
Furthermore, in the realm of time series anomaly detection, in the recent model called ContextDA [13], the discriminator aligns source/target domain windows without leveraging label information in the source domain, which makes the alignment ineffective. Our model leverages source labels and anomaly injection for better feature extraction, enhancing the alignment of normal samples. This is particularly vital due to the challenges in aligning anomalies across domains with different anomalous classes.
@@
In our study, we introduce Domain Adaptation Contrastive learning model for Anomaly Detection in time series (DACAD), a unique framework for UDA in multivariate time series anomaly detection, which leverages contrastive learning (CL) and focuses on contextual representations. It forms positive pairs based on proximity and negative pairs using anomaly injection, subsequently learning their embeddings with both supervised CL in the source domain and self-supervised CL in the target domain. To sum up, we make the following contributions:
@@
  • We introduce DACAD, a pioneering CL framework for multivariate time series anomaly detection with UDA. It is a novel deep UDA model based on contrastive representation utilizing a labeled source dataset and synthetic anomaly injection to overcome the lack of labeled data in TSAD, distinguishing between normal and anomalous patterns effectively.
@@
  • Extending on the deep one-class classifier (DeepSVDD) [14], our proposed one-class classifier, the CEC, operates on the principle of spatial separation in the feature space by leveraging the existing label information in the source domain, aiming to bring “normal” sample representations closer to the center and distancing anomalous ones. This spatial adjustment is measured using a distance metric, establishing the basis for anomaly detection.
@@
  • Our comprehensive evaluation with real-world datasets highlights DACAD’s efficiency. In comparison with the recent TSAD deep models and UDA models for time series classification and UDA for anomaly detection, DACAD demonstrates superior performance, emphasizing the considerable importance of our study.
II Related Works
@@
In this section, we provide a brief survey of the existing literature that intersects with our research. We concentrate on two areas: UDA for time series and deep learning for TSAD.
@@
UDA for time series: In the realm of UDA for time series analysis, the combination of domain adaptation techniques with the unique properties of time series data presents both challenges and opportunities [15]. Traditional approaches such as Maximum Mean Discrepancy (MMD) [16] and adversarial learning frameworks [17] are geared towards minimizing domain discrepancies by developing domain-invariant features. These methods are vital in applications spanning medical, industrial, and speech data, with significant advancements in areas like sleep classification [18], energy safety detection [19], arrhythmia classification [20], and various forms of anomaly detection [13], fault diagnosis [21], and lifetime prediction [22].
@@
Time series specific UDA approaches like variational recurrent adversarial deep domain adaptation (VRADA) [7], pioneered UDA for multivariate time series, utilizing adversarial learning with an LSTM networks [23] and variational RNN [24] feature extractor. Convolutional deep domain adaptation for time series (CoDATS) [10], built on VRADA’s adversarial training but employed a convolutional neural network as the feature extractor. A metric-based method, time series sparse associative structure alignment (TS-SASA) [8] aligns intra and inter-variable attention mechanisms between domains using MMD. Adding to these, the CL for UDA of time series (CLUDA) [9] model offers a novel approach, enhancing domain adaptation capabilities in time series. All these methods share a common objective of aligning features across domains, each contributing unique strategies to the complex challenge of domain adaptation in time series classification data. However, they are ineffective when applied to TSAD tasks. Additionally, ContextDA [13], is a TSAD model that applies deep reinforcement learning to optimize domain adaptation, framing context sampling as a Markov decision process. However, it is ineffective when the anomaly classes change between the source and the target.
Recently, Large Language Models (LLMs) and Pretrained Models (PMs) have gained attention for addressing Unsupervised Domain Adaptation (UDA) in time series tasks [25, 26]. These models leverage their robust transfer learning capabilities and the benefits of extensive pre-training on a vast array of sequential data across different domains. This pre-training phase equips the models with comprehensive knowledge and advanced pattern recognition skills, allowing them to effectively adapt to and perform well on specific time series tasks through a relatively straightforward fine-tuning process [27, 28]. This extensive pre-training enables LLMs and PMs to adeptly handle time series anomaly detection, even when these tasks differ significantly from the domains on which they were originally trained. Additionally, diffusion model-based UDA methods have emerged as a promising approach, utilizing the inherent strengths of diffusion models in capturing dynamic changes over time [29]. By integrating a small amount of target domain data to guide the adaptation process, these models can achieve effective domain transfer [30, 31]. However, both LLMs and diffusion models are not without their challenges. Issues such as computational efficiency and training stability are prevalent and continue to be significant areas of research and discussion in the field [29, 28].
@@
Deep Learning for TSAD: The field of Time Series Anomaly Detection (TSAD) has advanced significantly, embracing a variety of methods ranging from basic statistical approaches to sophisticated deep learning techniques [32, 33, 34]. Notably, deep learning has emerged as a promising approach due to its autonomous feature extraction capabilities [35]. TSAD primarily focuses on unsupervised [36, 37, 38, 39] and semi-supervised methods [40, 41, 42, 43] to tackle the challenge of limited labelled data availability. Unsupervised methods like OmniAnomaly [44], GDN [45], and BeatGAN [46] are especially valuable in scenarios with sparse anomaly labels, whereas semi-supervised methods leverage available labels effectively.
@@
Advanced models such as LSTM-NDT [37] and THOC [47] excel in minimizing forecasting errors and capturing temporal dependencies. However, models like USAD [38] face challenges with long time series data due to error accumulation in decoding. Additionally, unsupervised representation learning, exemplified by SPIRAL [48] and TST [49], shows impressive performance, albeit with scalability issues in long time series. Newer models like TNC [50] and T-Loss [51] aim to overcome these challenges using methods such as time-based negative sampling.
@@
Furthermore, contrastive representation learning, crucial in TSAD for pattern recognition, groups similar samples together while distancing dissimilar ones [52]. It has been effectively employed in TS2Vec [36] for multi-level semantic representation and in DCdetector [53], which uses a dual attention asymmetric design for permutation invariant representations. Recently, PatchAD [54] designs a lightweight patch-based MLP-mixer based on contrastive learning for TSAD.
In recent years, several works have been proposed that design universal models for various time series tasks, including anomaly detection. For example, TimesNet [55] transforms 1D time series data into 2D tensors to better capture and model complex temporal variations through multi-periodicity, and it improves performance across various time series analysis tasks. Additionally, some methods based on large models or pre-trained models have been proposed. For example, Zhou et al. [25] leverage pre-trained models, preserving their original self-attention and feedforward layers, to achieve state-of-the-art performance in various time series domain tasks by fine-tuning, demonstrating their effectiveness and universality. There are also specific methods for TSAD using large models or pre-trained models, such as AnomalyLLM [56] and AnomalyBERT [57].
III DACAD
@@
Problem formulation: Given an unlabeled time series dataset $T$ (target), the problem is to detect anomalies in $T$ using a labeled time series dataset $S$ (source) from a related domain.
@@
In this section, we present DACAD, which uses temporal correlations ingeniously and adapts to differences between source and target domains. It starts with a labeled dataset in the source domain, featuring both normal and anomalous instances. Both real and synthetic (injected) anomalies aid in domain adaptation, thereby training in a manner that improves generalizability on a wide range of anomaly classes.
@@
DACAD’s core is CL as inspired by a recent UDA for time series work [9], which strengthens its ability to handle changes between domains by improving the feature representation learning in source and target domains. As described in subsection III-C, in the target domain, we use a self-supervised contrastive loss [58] by forming triplets to minimize the distance between similar samples and maximize the distance between different samples. Additionally, in the source domain, we leverage label information of both anomaly and normal classes and propose an effective supervised contrastive loss [59] named supervised mean-margin contrastive loss. A Temporal Convolutional Network (TCN) in DACAD captures temporal dependencies, generating domain-invariant features. A discriminator ensures these features are domain-agnostic, leading to consistent anomaly detection across domains. DACAD’s strength lies in its dual ability to adapt to new data distributions and to distinguish between normal and anomalous patterns effectively. In DACAD, time series data is split into overlapping windows of size $WS$ with a stride of 1, forming detailed source ($S$) and target ($T$) datasets. Windows in the source domain are classified based on anomaly presence: those with anomalies are marked as anomalous windows ($S_{\text{anom}}$), and others as normal ($S_{\text{norm}}$). Fig. 1 shows the DACAD architecture, and Algorithm 1 details its steps. The following subsections explore DACAD’s components and their functionalities.
@@
5:Create $S_{\text{inj}}$, $T_{\text{inj}}$triangleright Anomaly Injection (III-A)
6:Form $S_{\text{triplets}}$ and $T_{\text{triplets}}$triangleright Pair Selection (III-B)
@@
8:    Compute $\phi^{R}(S)$,$\phi^{R}(T)$,$\phi^{R}(S_{\text{Triplets}})$,$\phi^{R}(T_{\text{Triplets}})$triangleright (III-C)
@@
11:    Compute $\mathcal{L}_{\text{Disc}}$ using Eq. (3), $\phi^{R}(S)$ and $\phi^{R}(T)$triangleright (III-D)
12:    Compute $\mathcal{L}_{\text{Cls}}$ using Eq. (III-E) and $\phi^{R}(S)$triangleright (III-E)
@@
III-A Anomaly Injection
In the anomaly injection phase, we augment the original time series windows through a process of negative augmentation, thereby generating synthetic anomalous time series windows. This step is applied to all windows in the target domain ($T$) and all normal windows in the source domain ($S_{\text{norm}}$). We employ the anomaly injection method outlined in [60], which encompasses five distinct types of anomalies: Global, Seasonal, Trend, Shapelet, and Contextual. This procedure results in the creation of two new sets of time series windows. The first set, $S_{\text{inj}}$, consists of anomaly-injected windows derived from the normal samples in the source domain. The second set, $T_{\text{inj}}$, comprises anomaly-injected windows originating from the target domain data.
III-B Pair Selection
@@
III-C Representation Layer ($\phi^{R}$)
In our model, we employ a TCN [61] for the representation layer, which is adept at handling both multivariate time series windows. This choice is motivated by the TCN’s ability to capture temporal dependencies effectively, a critical aspect in time series analysis. The inputs to the TCN are the datasets and triplets from both domains, specifically $S$, $T$, $S_{\text{Triplets}}$, and $T_{\text{Triplets}}$. The outputs of the TCN, representing the transformed feature space, are